# Complex variational autoencoders admit Kähler structure

### v1

@@ 5.3 A log-likelihood sum potential induces a metric asymptotically close to the Fisher information metric
$\partial\overline{\partial}\Psi_{i}=J^{\dagger}(z)\Sigma_{i}^{-1}J(z)=J^{\dagger}(z)\Gamma^{\dagger}\Gamma J(z)\succeq 0,$ (98)
@@
Proof. By Lemma 1, it follows
$\mathbb{E}_{w}[\partial_{\alpha}\partial_{\overline{\beta}}\Psi]\approx(\partial_{\alpha}\mu)^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\mu).$ (101)
@@
$\partial_{\alpha}\Psi_{i}$ $=(\partial_{\alpha}v)^{\dagger}\Sigma_{i}^{-1}v+v^{\dagger}\Sigma_{i}^{-1}(\partial_{\alpha}v)$ (103)
$\approx(M\Sigma_{i}^{-1}v)^{\dagger}\Sigma_{i}^{-1}v+v^{\dagger}\Sigma_{i}^{-1}(M\Sigma_{i}^{-1}v)$ (104)
$=v^{\dagger}(\Sigma_{i}^{-1}M^{\dagger}\Sigma_{i}^{-1}+\Sigma_{i}^{-1}M\Sigma_{i}^{-1})v=v^{\dagger}Pv.$ (105)
Again, there exists a weight, or weights, sufficiently large so that one, or multiple, $i\in I$ dominate. It is the case that $\Sigma_{i\in I}^{-1}\approx\Sigma(z)$. Thus,
@@ 5.4 Inference
It is highly distinguishable that sampling according the intrinsic decoder geometry is a less meaningful task than sampling according to the encoder geometry as in chadebec2022geometricperspectivevariationalautoencoders. The decoder metric does not characterize the probability distributions of the model with respect to generative inference. The decoder geometry measures latent perturbations into ambient pixel-space changes. It captures geometry of representations, not geometry of inference.
When working with our metrics, we will generally consider admissible metrics closely related to 27 and our metric sufficiently close (up to constants)
@@
thus our regularization has an enhanced reformulation
@@ Appendix A Our Kähler potential and the Fisher metric analytic equivalence
$\ \ \ \ \ \ \ \ \ \ \ \ -\mathrm{Tr}(\Sigma^{-1}\partial_{\overline{\beta}}\Sigma)\mathbb{E}\Big[v^{\dagger}\Sigma^{-1}(\partial_{\alpha}\Sigma)\Sigma^{-1}v\Big)\Big]+\mathbb{E}\Big[v^{\dagger}\Sigma^{-1}(\partial_{\alpha}\Sigma)\Sigma^{-1}vv^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\Sigma)\Sigma^{-1}v\Big)\Big]\Bigg)$ (138)
$=\rho^{4}\Bigg(2\text{Re}(\partial_{\alpha}\mu)^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\mu)-\mathrm{Tr}(\Sigma^{-1}\partial_{\alpha}\Sigma)\mathrm{Tr}(\Sigma^{-1}\partial_{\overline{\beta}}\Sigma)+\mathbb{E}\Big[v^{\dagger}\Sigma^{-1}(\partial_{\alpha}\Sigma)\Sigma^{-1}vv^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\Sigma)\Sigma^{-1}v\Big)\Big]\Bigg)$ (139)
@@
Let us illustrate more thoroughly that our regime in 5.3 is consistent with what we propose in this section. Let us assume parameter closeness and $\gamma I\preceq\Sigma_{i}$. Consider the following
@@
Appendix B Additional figures
Figure 5: We plot 100 randomly sampled diagonal entries $\partial_{\alpha}\partial_{\overline{\alpha}}\text{decoder}(z)$ in a histogram. The purpose of this plot is to convey the assumption $\partial_{\alpha}\partial_{\overline{\beta}}\mu(z)\approx 0$ is reasonable.

### v3

@@ 5.2 The Fisher information metric and the Hessian of the KL divergence as a Kähler potential
$\partial\overline{\partial}\Psi_{i}=J_{z}^{\dagger}(z)\Sigma_{i}^{-1}J_{z}(z)+(J_{\overline{z}}^{\dagger}(z)\Sigma_{i}^{-1}J_{\overline{z}}(z))^{T}\succeq 0,$ (103)
@@
Remark (non-rigorous, no proof provided). It can be noted that the potentials $K_{1}=\rho^{2}\log(\sum_{i}\exp\{-\Psi_{i}/\rho^{2}\})$, $K_{2}=-\rho^{2}\log(\sum_{i}\exp\{-\Psi_{i}/\rho^{2}\})$ are not (necessarily) plurisubharmonic.
@@
Proof. By Lemma 1, it follows for some $j$
$\mathbb{E}_{w}[\partial_{\alpha}\partial_{\overline{\beta}}\Psi]\approx(\partial_{\alpha}\mu_{j})^{\dagger}\Sigma_{j}^{-1}(\partial_{\overline{\beta}}\mu_{j})+(\partial_{\overline{\beta}}\mu_{j})^{\dagger}\Sigma_{j}^{-1}(\partial_{\alpha}\mu_{j}).$ (106)
@@
$\partial_{\alpha}\Psi_{i}$ $=(\partial_{\alpha}v)^{\dagger}\Sigma_{j}^{-1}v+v^{\dagger}\Sigma_{j}^{-1}(\partial_{\alpha}v)$ (108)
$\approx(M\Sigma_{j}^{-1}v)^{\dagger}\Sigma_{j}^{-1}v+v^{\dagger}\Sigma_{j}^{-1}(M\Sigma_{j}^{-1}v)$ (109)
$=v^{\dagger}(\Sigma_{j}^{-1}M^{\dagger}\Sigma_{j}^{-1}+\Sigma_{j}^{-1}M\Sigma_{j}^{-1})v=v^{\dagger}Pv.$ (110)
Again, there exists a weight, or weights, sufficiently large so that one, or multiple, $j\in I$ dominate. It is the case that $\Sigma_{j\in I}^{-1}\approx\Sigma(z)$. Thus,
@@ 5.4 Inference
It is important to distinguish that sampling according the intrinsic decoder geometry is a less meaningful task than sampling according to the encoder geometry as in chadebec2022geometricperspectivevariationalautoencoders. The decoder metric does not characterize the probability distributions of the model with respect to generative inference. The decoder geometry measures latent perturbations into ambient pixel-space changes. It captures geometry of representations, not geometry of inference.
When working with our metrics, we will generally consider admissible metrics closely related to 28 and our metric sufficiently close (up to constants)
@@
thus our regularization has an enhanced reformulation with notation $A=iA_{i\overline{j}}dz^{i}\wedge d\overline{z}^{j}$
@@ Appendix A Our Kähler potential and the Fisher metric analytic equivalence
$\ \ \ \ \ \ \ \ \ \ \ \ -\mathrm{Tr}(\Sigma^{-1}\partial_{\overline{\beta}}\Sigma)\mathbb{E}\Big[v^{\dagger}\Sigma^{-1}(\partial_{\alpha}\Sigma)\Sigma^{-1}v\Big]+\mathbb{E}\Big[v^{\dagger}\Sigma^{-1}(\partial_{\alpha}\Sigma)\Sigma^{-1}vv^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\Sigma)\Sigma^{-1}v\Big]\Bigg)$ (143)
$=\rho^{4}\Bigg(2\text{Re}(\partial_{\alpha}\mu)^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\mu)-\mathrm{Tr}(\Sigma^{-1}\partial_{\alpha}\Sigma)\mathrm{Tr}(\Sigma^{-1}\partial_{\overline{\beta}}\Sigma)+\mathbb{E}\Big[v^{\dagger}\Sigma^{-1}(\partial_{\alpha}\Sigma)\Sigma^{-1}vv^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\Sigma)\Sigma^{-1}v\Big]\Bigg)$ (144)
@@
Let us examine the regime under the circumstance $\min_{i}\Psi_{i}$ is chosen. Let us assume parameter closeness and $\gamma I\preceq\Sigma_{i}$. With a slight lack of rigor, consider the following
@@
Appendix B Expectation and covariance relations
In this section, we draw connections between the primary section of Section 5.3 and Appendix A.
We previously had
$g_{\alpha\overline{\beta}}=\mathbb{E}_{x|z}[\partial_{\alpha}\partial_{\overline{\beta}}\Psi]+\frac{1}{\rho^{2}}\text{Cov}_{x|z}(\partial_{\alpha}\Psi,\partial_{\overline{\beta}}\Psi)$ (161)
$h_{\alpha\overline{\beta}}=\mathbb{E}_{w(i)}[\partial_{\alpha}\partial_{\overline{\beta}}\Psi_{i}]+\frac{1}{\rho^{2}}\text{Cov}_{w(i)}(\partial_{\alpha}\Psi_{i},\partial_{\overline{\beta}}\Psi_{i}).$ (162)
Let us take $i\sim w(i),x|z,i\sim p_{i}(\cdot|z)$. By total expectation,
$\mathbb{E}_{i,x|z}[\partial_{\alpha}\partial_{\overline{\beta}}\Psi(i,x;z)]$ $=\mathbb{E}_{i\sim w(i)}[\mathbb{E}_{x|z,i}[\partial_{\alpha}\partial_{\overline{\beta}}\Psi_{i}(x;z)]]$ (163)
$\text{Cov}_{i,x|z}(\partial_{\alpha}\Psi(i,x;z),\partial_{\overline{\beta}}\Psi(i,x;z))$ $=\mathbb{E}_{i\sim w(i)}[\text{Cov}_{x|z,i}(\partial_{\alpha}\Psi_{i}(x;z),\partial_{\overline{\beta}}\Psi_{i}(x;z))]$ (164)
$\ \ \ \ \ \ +\text{Cov}_{i\sim w(i)}(\mathbb{E}_{x|z,i}[\partial_{\alpha}\Psi_{i}(x;z)],\mathbb{E}_{x|z,i}[\partial_{\overline{\beta}}\Psi_{i}(x;z)]).$ (165)
Define the joint object
$G_{\alpha\overline{\beta}}(z)=\mathbb{E}_{i,x|z}[\partial_{\alpha}\partial_{\overline{\beta}}\Psi_{i}(x;z)]+\frac{1}{\rho^{2}}\text{Cov}_{i,x|z}(\partial_{\alpha}\Psi(i,x;z),\partial_{\overline{\beta}}\Psi(i,x;z)).$ (166)
Thus,
$G_{\alpha\overline{\beta}}(z)$ $=\mathbb{E}_{i\sim w(i)}\Bigg[\mathbb{E}_{x|z,i}[\partial_{\alpha}\partial_{\overline{\beta}}\Psi_{i}(x;z)]+\frac{1}{\rho^{2}}\text{Cov}_{x|z,i}(\partial_{\alpha}\Psi_{i}(x;z),\partial_{\overline{\beta}}\Psi_{i}(x;z))\Bigg]$ (167)
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\frac{1}{\rho^{2}}\text{Cov}_{i\sim w(i)}(\mathbb{E}_{x|z,i}[\partial_{\alpha}\Psi_{i}(x;z)],\mathbb{E}_{x|z,i}[\partial_{\overline{\beta}}\Psi_{i}(x;z)]).$ (168)
We conclude
$G_{\alpha\overline{\beta}}(z)$ $\propto\mathbb{E}_{w(i)}\Bigg[\rho^{2}\Big(4\text{Re}(\partial_{\alpha}\mu_{i})^{\dagger}\Sigma_{i}^{-1}(\partial_{\overline{\beta}}\mu_{i})+2\mathrm{Tr}\big(\Sigma_{i}^{-1}(\partial_{\alpha}\Sigma_{i})\Sigma_{i}^{-1}(\partial_{\overline{\beta}}\Sigma_{i})\big)\Big)\Bigg]$ (169)
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\frac{1}{\rho^{2}}\mathrm{Cov}_{w(i)}\Big(\mathbb{E}_{x|z,i}[\partial_{\alpha}\Psi_{i}],\mathbb{E}_{x|z,i}[\partial_{\overline{\beta}}\Psi_{i}]\Big).$ (170)
Appendix C Additional figures
Figure 6: We plot 100 randomly sampled diagonal entries $\partial_{\alpha}\partial_{\overline{\alpha}}(\text{decoder}(z).\texttt{mean}())$ in a histogram. The mean is approximately $0.10$ on MNIST data, so the pixel data is nontrivial in value compares to its second order Wirtinger derivatives. The purpose of this plot is to convey the assumption $\partial_{\alpha}\partial_{\overline{\beta}}\mu(z)\approx 0$ is reasonable. We plot the absolute value of each component, and the max across two elements $\alpha$.
Figure 7: We plot 100 randomly sampled diagonal entries $\partial_{\alpha}(\text{decoder}(z).\texttt{mean}())$ in a histogram. The mean is approximately $0.10$ on MNIST data, so the pixel data is nontrivial in value compares to its second order Wirtinger derivatives. The purpose of this plot is to convey the first derivatives are much larger in magnitude on average with our architectures than the mixed second derivatives. Moreover, the plot with respect to $\overline{\alpha}$ is the same, thus the decoder is not holomorphic. We plot the absolute value of each component, and the max across two elements $\alpha$. The x-scale is consistent with Figure 6. As a last remark, note that $\max|\partial_{\overline{z}}f-\overline{\partial_{z}f}|=0$, but $\max|\partial_{z}f-\partial_{\overline{z}}f|=0.0281$ on our real data.